An amount of R4 317,26 was borrowed on 5 May at a simple interest rate of 15% per year. The loan will be worth R4 500 on

1.     16 August.

2.     12 August.

3.     21 August.

4.     9 October.

5.     none of the above.


1
Expert's answer
2021-05-21T07:27:12-0400

simple interest rate",S=P(1+rt)"

"t=5may" till 16 Aug

date value =228-125=103days

"t=103\/365"

"S=p(1+rt)=4317,26(1+15/100\times 103/365)=R4500"

Therefore, the loan will be worth R4 500 on 16 August.